# The Buffer Mechanism for Multi-Step Information Reasoning in Language Models

### v2

@@ 4 Understanding Horizontal Thinking Strategy With Buffer Mechanism
The Chain of Thought (CoT) approach, as a representative of horizontal thinking strategy, has become a prevalent mode of reasoning adopted by current large models, with numerous studies indicating that CoT can significantly enhance logical reasoning capability (Wei et al., 2022; Kojima et al., 2022). In this section, we will demonstrate that a Transformer model utilizing CoT can achieve arbitrary-step multi-step reasoning with only two layers. As shown in Fig. 6, we illustrate how the Transformer model employs CoT for multi-step reasoning. In the previously mentioned vertical thinking scenario, the model needed to allocate a new buffer ${\bm{W}}^{vo(1)}$ for storing the intermediate result [b] and [c]. In contrast, with CoT, the model generates a new token to separately store the new intermediate information, effectively replacing the information [a] in the original buffer ${\bm{W}}^{vo(0)}$ and ${\bm{I}}$ (identity matrix, which can also be treated as a buffer). Thus, CoT significantly enhances the model’s information processing capability through buffer reuse.
Figure 6: Illustration of how the Transformer model employs horizontal thinking strategy for multi-step reasoning task. The red, blue and green portions indicate the information activated during the first, second and third reasoning step, respectively.
5 Method for Improving Transformer’s Data-Efficiency
@@
A more intuitive approach is to replace ${\bm{W}}^{qk(l)}$ and ${\bm{W}}^{vo(l)}$ with ${\bm{W}}^{qk(l)}+\alpha^{(l)}I$ and ${\bm{W}}^{vo(l)}+\beta^{(l)}I$, where $\{\alpha^{(l)}\}_{l=1}^{L}$ and $\{\beta^{(l)}\}_{l=1}^{L}$ are learnable parameters. This Identity Matrix-Based Algorithm (denoted as IMBA) was first proposed in Boix-Adsera et al. (2023) and has been validated from both theoretical and empirical perspectives for its role in facilitating the model’s learning of one-step reasoning data. However, in multi-step reasoning tasks, IMBA may lead to information interference. For instance, if ${\bm{W}}^{vo(l)}$ is replaced with ${\bm{W}}^{vo(l)}+\beta^{(l)}{\bm{I}}$, the storage representation of the two pieces of information transitions from [a]${\bm{W}}^{vo}$+[b] to [a]${\bm{W}}^{vo}$+($\beta^{(l)}$[a]+[b]), which introduces interference among the different pieces of information. Therefore, this approach may not be effective in enhancing the model’s ability to perform multi-step reasoning.
Based on the understanding of buffer mechanism, we propose a Random Matrix-Based Algorithm (RMBA), specifically by substituting ${\bm{W}}^{qk(l)}$ and ${\bm{W}}^{vo(l)}$ with ${\bm{W}}^{qk(l)}+\alpha^{(l)}{\bm{Z}}^{(l-1)}$ and ${\bm{W}}^{vo(l)}+\beta^{(l)}{\bm{Z}}^{(l)}$, where $\{\alpha^{(l)}\}_{l=1}^{L}$ and $\{\beta^{(l)}\}_{l=1}^{L}$ are learnable parameters and $\{{\bm{Z}}^{(l)}\}_{l=1}^{L}$ is a set of fixed random matrix following $N(0,1/d_{m})$. In this case, the information storage representation changes from [a]${\bm{W}}^{vo}$+[b] to [a]${\bm{W}}^{vo}$+[a]${\bm{Z}}$+[b], effectively creating a new buffer.
Fig. 7 illustrates the different results of the two methods. The baseline is a three-layer transformer model, which fails to learn a two-step reasoning task with only 30,000 samples. Under various hyperparameter settings, when $\alpha^{(l)}_{\text{ini}}\cdot\beta^{(l)}_{\text{ini}}\geq 0$, the RMBA algorithm significantly enhances the model’s generalization ability, while the IMBA shows no effect. This experiment validates our buffer mechanism understanding. This result also demonstrates that multi-step reasoning is not achieved by simply “stacking" multiple single-step reasonings. An algorithm that enhances single-step reasoning may not be applicable to multi-step reasoning. Therefore, investigating the mechanisms of multi-step reasoning is an important and meaningful topic.
@@
Figure 7: The impact of different learnable parameters’ initial values, $\alpha^{(l)}$ and $\beta^{(l)}$, on the model’s reasoning ability. When $\alpha^{(l)}_{\text{ini}}=0$ or $\alpha^{(l)}$ and $\beta^{(l)}$ have the same sign initially, the model’s reasoning capability can be enhanced. The solid lines represent the training accuracy, while the dashed lines denote the test accuracy. Each experiment was conducted with 5 random seeds (each seed corresponds to one line). Detailed settings can be found in Appendix E.
5.1 Validation of Algorithm Performance on PrOntoQA Dataset
PrOntoQA dataset (Saparov & He, 2022) is constructed to test the reasoning capability of language models. We use the data generation procedure to create a 0.5M dataset of 3-step reasoning question-answer pairs, formatted as shown in Fig. 8(a).
We use a 12-layer, 12-head GPT-2 small model for this dataset. For the IMBA, we followed the settings provided in the open-source code (Boix-Adsera et al., 2023), which includes 144 learnable parameters $\alpha^{(l,h)}$ and 144 learnable parameters $\beta^{(l,h)}$. For RMBA, we replace ${\bm{W}}^{qk(l,h)}$ and ${\bm{W}}^{vo(l,h)}$ with ${\bm{W}}^{qk(l,h)}+\alpha^{(l,h)}{\bm{Z}}^{(l-1,h)}$ and ${\bm{W}}^{vo(l,h)}+\beta^{(l,h)}{\bm{Z}}^{(l,h)}$. Fig. 8(b) illustrates the impact of different algorithms on the performance of the original GPT-2 small model (baseline) on this dataset. In this figure, accuracy reflects the model’s precision in predicting True or False for the last token. The results indicate that the model utilizing RMBA achieves the highest data efficiency, reducing the number of iterations required to attain generalization by 75% compared to the baseline model.
To verify whether these algorithms enable the model to learn the essence of the reasoning process, we inserted several random tokens at random positions in each test sentence and recorded the accuracy of the different models. Fig. 8(b) shows the accuracy of each model outputting the correct answer as the number of random tokens increases. The IMBA models exhibit the least stability, while the RMBA models maintain the same level of stability as the baseline.
Figure 8: (a) Example of PrOntoQA dataset. For each algorithm, we initialize the learnable parameter $\alpha^{(l)},\beta^{(l)}$ at 0.01 and 0.05. (b) Accuracy curve of GPT-2 trained with RMBA and IMBA. (c) Stability of different GPT-2 models.
6 Related Work
Language model reasoning There have been numerous experimental and theoretical studies on language model reasoning. Abbe et al. (2022) examines the reasoning capabilities of neural networks using the Pointer Value Retrieval (PVR) benchmark, which was originally introduced in Zhang et al. (2021). Sharma et al. (2023) demonstrates that applying low-rank approximation of the certain weight layers to themselves can enhance reasoning performance across various tasks, and Chen et al. (2024a) explains this phenomenon in a two-layer Transformers model. Wang et al. (2024a) investigate both ID and OOD reasoning abilities on two synthetic tasks involving composition and comparison. Jiang et al. (2024) reveals that the reasoning process in Large Language Models (LLMs) is influenced by token bias and that these models continue to face challenges when dealing with more complex logical tasks. Abbe et al. (2024) introduces the distribution locality and shows Transformers require low locality. Zhang et al. (2024b); Luo et al. (2021); Zhou et al. (2022) shows that small initialization can facilitate model reasoning. Brinkmann et al. (2024) investigates the mechanism by which language models output intermediate reasoning paths in multi-step reasoning tasks.
Understanding the mechanism of neural model Our work builds upon previous studies on the attention mechanism (Voita et al., 2019; Vig, 2019; Kovaleva et al., 2019; Kobayashi et al., 2020). Numerous researchers have proposed various approaches to identify the roles of different heads in Transformers (Vig et al., 2020; Jeoung & Diesner, 2022; Wang et al., 2022; Conmy et al., 2023; Merullo et al., 2023; Guo et al., 2023; Wang & Weinan, 2024; Amsel et al., 2024; Li et al., 2024; Wang et al., 2024c). These methods predominantly employ the concept of perturbation. Similar to the observations made by Wang et al. (2023) and Dutta et al. (2024), who noted that large language models typically perform information aggregation in shallow layers and information induction in deeper layers, we have also observed comparable phenomena in our study. The idea of symbolic datasets is inspired by Poli et al. (2024); Zhang et al. (2024c). There have also been some insightful theoretical works on feedforward neural networks. A series of studies have explored neural network preferences and generalization from the perspectives of regularization and frequency, etc. (Xu et al., 2019; Wang et al., 2024b; Jacot et al., 2018, 2020; Arora et al., 2019a, 2018). And Wu & Su (2023); Wang & Wu ; Arora et al. (2022); Li et al. (2021); Wu et al. (2018); Zhu et al. (2018); Arora et al. (2019b) investigates the dynamical behavior of neural networks, while Ren et al. (2024) examines the factors influencing neural network generalization.
In-context learning and induction head Our work primarily investigates the model’s ability to perform in-context learning. The concept of in-context learning (ICL) was first introduced by Brown et al. (2020). Since then, a series of studies have utilized induction heads to investigate ICL, yielding remarkable research outcomes (Olsson et al., 2022; Garg et al., 2022; Wang et al., 2022; Müller et al., 2021; Goldowsky-Dill et al., 2023; Bietti et al., 2024; Nichani et al., 2024; Edelman et al., 2024; Chen et al., 2024b; Todd et al., 2023; Chen & Zou, 2024). It is worth noting that induction heads can be considered as a special case of multi-step reasoning tasks with reasoning step equals to $1$. However, multi-step reasoning is not a simple linear combination of single-step reasoning. In our work, we study the mechanism that enables multi-step reasoning, which has not been explored in previous studies.
@@ 7 Discussion
Conclusion. In this study, we investigated the buffer mechanism employed by Transformer models when performing multi-step reasoning using the vertical thinking strategy and horizontal thinking strategy. When utilizing the vertical thinking strategy, the model stores different intermediate results in separate buffers and transfers the reasoning results with same-token matching. In contrast, when applying the horizontal thinking strategy, the model reuses the existing buffers to store intermediate results, enabling reasoning over an arbitrary number of steps. We validated that the buffer mechanism is a key factor in enabling the model’s ID and OOD generalization capabilities. Based on the buffer mechanism, we proposed a tailored approach, RMBA, to enhance the model’s multi-step reasoning ability, significantly improving data efficiency when training GPT-2 on PrOntoQA datasets.
@@
Limitations and Future Work. Our current work lacks in-depth theoretical analysis. In future work, we aim to conduct deeper theoretical modeling for multi-step reasoning problems and extend the buffer mechanism to other types of multi-step reasoning tasks.
Appendix A Experimental settings

### v3

@@ 4 Buffer Mechanism
Understanding HTS With Buffer. Numerous studies have shown that CoT, a canonical representative of HTS, can significantly enhance logical reasoning capability (Wei et al., 2022; Kojima et al., 2022). Fig. 3(b) presents a schematic diagram illustrating how the model implements this process. In the previously mentioned VTS scenario, the model needed to allocate new buffer ${\bm{W}}^{vo(1)}$, ${\bm{W}}^{vo(2)}$ for storing the intermediate result [b] and [c], respectively. In contrast, with CoT, the model generates a new token to separately store the new intermediate information, effectively replacing the information [a] in the original buffer ${\bm{W}}^{vo(1)}$, ${\bm{W}}^{vo(0)}$ and $\bm{I}$} (identity matrix, which can also be treated as a buffer). Thus, CoT performs multi-step reasoning tasks through buffer reuse.
Unlike the vertical thinking strategy, which requires alignment across multiple weight matrices, the horizontal thinking strategy only requires a few layers to satisfy $\text{Ker}^{(l)}\approx{\bm{I}}$. This significantly reduces the difficulty of model training. These architectural efficiencies potentially underlie the empirically observed enhancement of reasoning performance in large language models employing CoT methodologies.
5 Experiments
This section presents the experimental validation of the buffer theory. To demonstrate multi-step reasoning without introducing additional complexity, we utilized a 3-layer, single-head Transformer model to learn from a 2-step reasoning dataset. Specific Experimental settings can be found in Appendix B. After training, the Transformer exhibits generalization capability in both in-distribution and out-of-distribution data (Fig. 4).
Figure 4: Accuracy curve during VTS training. After training, the Transformer exhibits generalization capability in in-distribution (100% accuracy) and out-of-distribution data (82% accuracy).
Fig. 5 presents the experimental validation of Eq.(3)(4). We visualize the computed values of $h^{(i)}({\bm{X}}_{tgt})$ and $\text{Ker}^{(i)}$, $i=1,2$ with the model parameters after training. Notably, in the $\texttt{token}_{\texttt{OOD}}$ region, $\texttt{token}_{\texttt{OOD}}$ and $h^{(2)}({\bm{X}}_{tgt})$ exhibit a diagonal structure similar to an identity matrix $\bm{I}$}, ensuring the model’s OOD generalization capability.
To establish a strong correlation between the model’s OOD generalization and its ability to use the buffer mechanism for same-token matching, we define a metric for this capability, the Matching Score (MS):
$\text{MS}(h^{{(l)}})=\mathbb{E}_{{\bm{X}}}[\text{Trace}(\sigma(h^{(l)}({\bm{X}})))]/n,$ (5)
where ${\bm{X}}\in\mathbb{R}^{n\times d_{m}}$ is sampled from $\texttt{token}_{\texttt{ID}}$ or $\texttt{token}_{\texttt{OOD}}$ for expectation $\mathbb{E}_{{\bm{X}}}$, and $\sigma$ takes the softmax operation. The level of diagonalization of $\text{Ker}^{(l)}$ serves as the intrinsic driver for achieving same-token matching; thus, we also define the following Kernel Score (KS):
$\color[rgb]{0,0,0}\text{KS}(\text{Ker}^{(l)})=\text{Trace}(\sigma(\text{Ker}^{(l)}))/d_{m}$}. (6)
Figure 5: Heatmap of $h^{(1)}({\bm{X}}_{tgt})$, $h^{(2)}({\bm{X}}_{tgt})$, $\text{Ker}^{(1)}$ and $\text{Ker}^{(2)}$. According to (c)(d), and Eq.(4), the diagonal structure of the kernel matrix induces a diagonal structure in the matching matrix, even when ${\bm{X}}_{\text{tgt}}$ is sampled from the $\texttt{token}_{\texttt{OOD}}$.
Fig. 4 and Fig. 6 illustrates the accuracy curve and the dynamic changes in the matching score and kernel score during training. It is observed that the increase in the model’s ID and OOD generalization coincide with the increases in the model’s matching score and kernel score. By computing the cosine similarity between different buffers (${\bm{W}}^{vo(0)},{\bm{W}}^{vo(1)},{\bm{W}}^{vo(2)}$ and $I$), we observe that these buffers are nearly pairwise orthogonal(Appendix D). Based on the above experimental analysis, we conclude that the Transformer model indeed employs the buffer mechanism for vertical thinking.
Figure 6: (a) The dynamic evolution of the model’s matching score. The red and blue lines represent the matching scores for the first and second layers, respectively. Solid and dashed lines indicate the matching scores when $\bm{X}$} is drawn from $\texttt{token}_{\texttt{ID}}$ and $\texttt{token}_{\texttt{OOD}}$, respectively. (b) The kernel scores for the first (red) and second (blue) layers.
According to the Buffer mechanism, in theory, by setting the model’s weights in the following way, an $L$-layer model can achieve (L1)(L-1)-step reasoning without the extra training process:
${\bm{W}}^{q(0)}={\bm{W}}^{q(1)}=I,\ {\bm{W}}^{q(l)}={\bm{W}}^{vo(l-1),\mathsf{T}},\ l\geq 2,$
${\bm{W}}^{k(0)}=\sum_{i=1}^{[l_{\text{seq}}/2]}{\bm{p}}_{2i}{\bm{p}}_{2i-1}^{\mathsf{T}},\ {\bm{W}}^{k(l)}={\bm{W}}^{vo(0),\mathsf{T}},\ l\geq 1,$
where $\{{\bm{W}}^{vo(l)}\}_{l=1}^{L}$ are set as random matrices and the projection layer satisfies ${\bm{W}}^{p}={\bm{W}}^{vo(L),\mathsf{T}}{\bm{W}}^{\text{emb},\mathsf{T}}$. Experimental validation can be found in Appendix D.
Real World LLMs. In Appendix G, we provide additional definitions for methods to compute the matching score and kernel score in multi-head models and perform these calculations in the real world LLMs such as GPT, Phi, Llama, and Qwen. Similar phenomena, such as weight alignment, provide evidence for the presence of the buffer mechanism in real language models.
Horizontal Thinking Experiments. For HTS, we will demonstrate that a Transformer model utilizing CoT can achieve arbitrary multi-step reasoning with only 2 layers. We trained a 2-layer Transformer with the 13-length single-step reasoning data. During the testing phase, we fed the model’s output back into the model. Through this CoT process, the model can perform 2-step, 3-step, or even higher-step reasoning and it can also generalize to sentence lengths beyond the 13th position. Fig. 7 shows the relationship between the number of reasoning steps and CoT accuracy. Our 2-layer model is able to maintain an accuracy of over 57.6% even when performing complex 6-step reasoning tasks. The specific information flow can be found in Appendix F.
Figure 7: The relationship between the number of reasoning steps and CoT accuracy. The CoT accuracy curve represents the performance when all reasoning steps are predicted correctly using CoT, while the CoT last token accuracy curve records the performance based solely on whether the final reasoning step is correct when the previous correct reasoning result is given.
6 Improving Transformer’s Data-Efficiency
@@
[…]
In RMBA, when $\alpha^{(l)}\cdot\beta^{(l)}\geq 0$, the diagonal elements of $({\bm{W}}^{vo{(l-1)},\mathsf{T}}+\beta^{(l-1)}{\bm{Z}}^{(l-1),\mathsf{T}})({\bm{W}}^{qk{(l)}}+\alpha^{(l)}{\bm{Z}}^{(l-1)})$ will exhibit positive increments, thereby enhancing the model’s reasoning generalization capability. Fig. 8 supports this perspective, while also showing that IMBA, based on induction heads, fails to improve multi-step reasoning performance. To strengthen the credibility of our results, we conducted a comprehensive sweep of hyperparameters such as weight decay, learning rate, and hidden dimensions ($d_{m}$ and $d_{k}$). The findings indicate that RMBA parameterization can more robustly reach a high accuracy by a certain number of training steps across a wider range of hyperparameters. Detailed experimental results can be found in Appendix E.
These results also demonstrate that multi-step reasoning is not achieved by simply “stacking" multiple single-step reasonings. An algorithm that enhances single-step reasoning may not be applicable to multi-step reasoning. Therefore, investigating the multi-step reasoning is an important and meaningful topic.
@@
Figure 8: The accuracy comparison for the Baseline model, RMBA model, and IMBA model. The baseline is a three-layer transformer model, which fails to learn a two-step reasoning task with only 30,000 samples. The solid lines represent the training accuracy, while the dashed lines denote the test accuracy. For both RMBA and IMBA models, we set 9 different initialization parameter values $\alpha^{(l)}_{\text{ini}}$ and $\beta^{(l)}_{\text{ini}}$ and each experiment was conducted with 5 random seeds (each seed corresponds to one line). When $\alpha^{(l)}_{\text{ini}}=0$ or $\alpha^{(l)}_{\text{ini}}\cdot\beta^{(l)}_{\text{ini}}>0$, the model’s reasoning capability can be enhanced. Detailed settings can be found in Appendix E.
Validation of Algorithm Performance on Multiple Real Multi-Step Reasoning Datasets. In this section, we evaluate the performance of our algorithm using published real-world datasets. The datasets used include Clutrr, RuleTaker, StepGame, LogicInference, LogicAsker, PararulePlus, and PrOntoQA. These datasets assess the model’s multi-step reasoning capabilities from different perspectives and scenarios. Examples of the datasets and their corresponding training curves are provided in Appendix E.
The 12-layer GPT-2 model is employed for this task. For RMBA, we replace ${\bm{W}}^{qk(l,h)}$ and ${\bm{W}}^{vo(l,h)}$ with ${\bm{W}}^{qk(l,h)}+\alpha^{(l,h)}{\bm{Z}}^{(l-1,h)}$ and ${\bm{W}}^{vo(l,h)}+\beta^{(l,h)}{\bm{Z}}^{(l,h)}$, respectively. To demonstrate the effectiveness of our method, we train each dataset for 5 epochs using a learning rate of $1e\text{-}5$. The additional hyperparameters introduced in RMBA are initialized to 0.05. Training results are summarized in Table 1. For tasks including Clutrr, RuleTalker, StepGame, and LogicInference, RMBA significantly improves model performance. For relatively simpler tasks such as LogicAsker, PararulePlus, and PrOntoQA, although both the base model and RMBA achieve nearly 100% accuracy, the training curves presented in Fig. 16 (Appendix E) show differences in the cost required to reach generalization capability. Table 2 reports the number of epochs required to reach 95% accuracy. As shown, RMBA reduces computational costs by more than 65% compared to the base model.
Table 1: Final-State Accuracy (%) Comparison. The training curves for all datasets are shown in Fig. 16 (Appendix E).
Dataset Baseline RMBA (ours) IMBA
Clutrr 34.3 42.8 42.6
RuleTaker 67.0 83.5 74.9
StepGame 35.9 38.9 35.3
LogicInference 68.1 78.3 49.2
LogicAsker 99.7 100.0 97.1
PararulePlus 99.8 99.8 99.7
PrOntoQA 100.0 100.0 100.0
Table 2: Epochs required to reach 95% accuracy and cost savings comparison
Dataset Baseline RMBA(ours) IMBA
LogicAsker 3.16 0.88 3.32
PararulePlus 2.71 0.73 2.02
PrOntoQA 0.95 0.32 1.25
@@ 7 Discussion
In this study, we investigated the vertical and horizontal thinking strategy employed by Transformer in a symbolic multi-step reasoning task from the perspective of the buffer mechanism. When utilizing the vertical thinking strategy, the model stores different intermediate results in separate buffers and transfers the reasoning results with the same-token matching. In contrast, when applying the horizontal thinking strategy, the model reuses the existing buffers to store intermediate results. We validated that the buffer mechanism is a key factor in enabling the model’s ID and OOD generalization capabilities. Based on the buffer mechanism, we proposed a tailored approach, RMBA, to enhance the model’s multi-step reasoning ability, significantly improving data efficiency when training GPT-2 on 7 reasoning datasets.
8 Limitation
@@
Our current work lacks an in-depth theoretical analysis. In future work, we aim to conduct deeper theoretical modeling for multi-step reasoning problems and extend the buffer mechanism to other types of multi-step reasoning tasks.
Appendix A Related Work
Language model reasoning There have been numerous experimental and theoretical studies on language model reasoning. Abbe et al. (2022) examines the reasoning capabilities of neural networks using the Pointer Value Retrieval (PVR) benchmark, which was originally introduced in Zhang et al. (2021). Sharma et al. (2023) demonstrates that applying low-rank approximation of certain weight layers to themselves can enhance reasoning performance across various tasks, and Chen et al. (2024a) explains this phenomenon in a two-layer Transformer model. Wang et al. (2024a) investigates both ID and OOD reasoning abilities on two synthetic tasks involving composition and comparison. Jiang et al. (2024) reveals that the reasoning process in Large Language Models (LLMs) is influenced by token bias and that these models continue to face challenges when dealing with more complex logical tasks. Abbe et al. (2024) introduces the distribution locality and shows that Transformers require low locality. Zhang et al. (2024b); Luo et al. (2021); Zhou et al. (2022) shows that small initialization can facilitate model reasoning. Brinkmann et al. (2024) investigates the mechanism by which language models output intermediate reasoning paths in multi-step reasoning tasks. Aubry et al. (2024) uncover a transformer block coupling phenomenon in a variety of LLMs by tracing the trajectories of individual tokens as they pass through transformer blocks. Recently, the multi-hop reasoning abilities of models have garnered attention in the LLM field. Kil et al. (2024); Li et al. (2024a) promote the use of Chain-of-Thought (CoT) reasoning by models to perform multi-step thinking through appropriate prompt engineering. Zhang et al. (2023) proposes a new retrieval framework for multiple reasoning paths. Dar et al. (2022); Li et al. (2024b); Yang et al. (2024); Shalev et al. (2024) pinpoint where models execute multi-step reasoning by causal intervention and observing neuron activation. Biran et al. (2024) introduces a method to enhance models’ reasoning abilities by repeatedly invoking intermediate layers. These works, especially the causal intervention experiments, have inspired our research. However, existing studies primarily conduct macro-level statistical analyses on actual complex large language models. While causal intervention methods can help us identify critical paths, our work builds upon this foundation by further exploring how models leverage their own weights to generate these key paths. Experiments on our symbolic datasets also facilitate more in-depth experimental and theoretical investigations.
Understanding the mechanism of neural model Our work builds upon previous studies on the attention mechanism (Voita et al., 2019; Vig, 2019; Kovaleva et al., 2019; Kobayashi et al., 2020). Numerous researchers have proposed various approaches to identify the roles of different heads in Transformers (Vig et al., 2020; Jeoung and Diesner, 2022; Wang et al., 2022; Conmy et al., 2023; Merullo et al., 2023; Guo et al., 2023; Wang and Weinan, 2024; Amsel et al., 2024; Li et al., 2024c; Wang et al., 2024c). These methods predominantly employ the concept of perturbation. Similar to the observations made by Wang et al. (2023) and Dutta et al. (2024), who noted that large language models typically perform information aggregation in shallow layers and information induction in deeper layers, we have also observed comparable phenomena in our study. The idea of symbolic datasets is inspired by Poli et al. (2024); Zhang et al. (2024c). There have also been some insightful theoretical works on feedforward neural networks. A series of studies have explored neural network preferences and generalization from the perspectives of regularization and frequency, etc. (Xu et al., 2019; Wang et al., 2024b; Jacot et al., 2018, 2020; Arora et al., 2019a, 2018). And Wu and Su (2023); Wang and Wu (2023); Arora et al. (2022); Li et al. (2021); Wu et al. (2018); Zhu et al. (2018); Arora et al. (2019b) investigates the dynamical behavior of neural networks, while Ren et al. (2024) examines the factors influencing neural network generalization.
In-context learning and induction head Our work primarily investigates the model’s ability to perform in-context learning. The concept of in-context learning (ICL) was first introduced by Brown et al. (2020). Since then, a series of studies have utilized induction heads to investigate ICL, yielding remarkable research outcomes (Olsson et al., 2022; Garg et al., 2022; Wang et al., 2022; Müller et al., 2021; Goldowsky-Dill et al., 2023; Bietti et al., 2024; Nichani et al., 2024; Edelman et al., 2024; Chen et al., 2024b; Todd et al., 2023; Chen and Zou, 2024). It is worth noting that induction heads can be considered as a special case of multi-step reasoning tasks with the reasoning step equal to $1$. However, multi-step reasoning is not a simple linear combination of single-step reasoning. In our work, we study the mechanism that enables multi-step reasoning, which has not been explored in previous studies.
Appendix B Experimental settings